Intro Graphics March 21, 2006 G. Farin

Programming Assignment # 3, tentative due: 4-10 percent of your your grade: 12.5

This project is about creating and displaying a 3D object.

Using either Loop subdivision or surfaces of revolution, implement displaying your object •in wireframe mode (that’s not very new!)

• smooth shaded using one, two, or three different point light sources • using different surface characteristics (have three available, such as two different colors, matte, and highly reflective)

• rotating your object with the light sources staying in place.